# On estimation and order selection for multivariate extremes via clustering

### v1

@@ 3.2.1 Some deterministic estimates
For $\epsilon,\delta>0$ as in Assumption 2, choose them small enough such that (30) is satisfied. Define the event
$E_{n}(\epsilon,\delta)=\{|W_{n}\cap B(\mathbf{a}_{i},\epsilon)|\geq|W_{n}|(p_{i}-\delta)\text{ for all }i\in\left\{1,\ldots,k\right\}\}.$ (33)
By Proposition 2.4 and the choice $\epsilon<r_{A}$, we have each $|W_{n}\cap B(\mathbf{a}_{i},\epsilon)|/|W_{n}|$ converges almost surely to $p_{i}$. Hence, with probability $1$, the event $E_{n}(\epsilon,\delta)$ happens eventually as $n\rightarrow\infty$, namely, $\Pr\left(\liminf_{n}\mathbf{1}\left\{E_{n}(\epsilon,\delta)\right\}=1\right)=1$. So, by Lemmas 3.2 and 3.5, for almost every outcome $\omega$ in the sample space $\Omega$, when $n$ is sufficiently large, we have when $m<k$ that
@@
The desired results follow if one takes $\limsup_{n}$ and $\liminf_{n}$ respectively in the two inequalities above, and then lets $\delta,\epsilon\rightarrow 0$. ∎
@@
Suppose $t>0$. If $m>k$, we have almost surely
@@
$\limsup_{n}P_{t}(W_{n};A_{k,n},\mathfrak{C}_{k,n})\leq 1-\left(r_{A}kp_{\min}\right)^{t}.$
@@
Putting together Propositions 3.6 and 3.7, and using the facts that $\bar{S}\in[0,1]$ and $P_{t}\in[0,1]$, we have almost surely that
@@ 4 Large deviation analysis of clustering-based spectral estimation
First, we pepare a Chernoff-Hoeffding-type bound for the sum of a Binomial random number of Bernoulli random variables, which may be of some independent interest.
@@
We only prove the (34) and the proof of (35) is similar. It follows from a version of Hoeffding’s inequality for Binomial [15, Equation (2.1)] that for any $m\geq 0$,
@@
where in the last inequality we have used the inequality $x+1\leq\exp\left(x\right)$, $x\in\mathbb{R}$. To obtain the second inequality in (34), it suffices to note that in view of [15, Equation (2.3)] one has $\mathcal{D}(q_{1}-r|q_{1})\geq 2r^{2}$. ∎
@@
Let $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$ be as defined in (18) and $\mathbf{a}_{\pi_{n}(i),n}^{k},p_{\pi_{n}(i),n}^{k}$ are as in Corollary 2.5. Note that an accurate estimation can be interpreted as that for small $x,y>0$, there exists permutation $\pi$, such that $D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})<x$ and $|p_{\pi(i),n}^{k}-p_{i}|<y$ for all $i\in\{1,\ldots,k\}$. Now consider the complement “large deviation” event
$E(x,y)=\bigcap_{\pi}\bigcup_{i=1}^{k}\left\{|\mathbf{a}_{\pi(i),n}^{k}-\mathbf{a}_{i}|>x\right\}\cup\left\{|p_{\pi(i),n}^{k}-p_{i}|>y\right\}.$
We have the following result.
@@
For any $x,y>0$,
@@
If $H_{n}\left(B(\mathbf{a}_{i},\epsilon)\right)=|W_{n}\cap B(\mathbf{a}_{i},\epsilon)|/|W_{n}|\geq p_{i}-\delta$ for all $i\in\{1,\ldots,k\}$, by Lemmas 3.3 and 3.5, as long as (30) holds, there exists a permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})<\epsilon^{\prime}$ and $|p_{\pi(i),n}^{k}-p_{i}|\leq c_{k}\delta$ for all $i\in\{1,\ldots,k\}$. Hence under (30), whenever $\epsilon^{\prime}\leq x$ or $c_{k}\delta\leq y$,
$\Pr\left(E(x,y)\right)\leq\Pr\left(\bigcap_{\pi}\bigcup_{i=1}^{k}\left\{D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})>\epsilon^{\prime}\right\}\cup\left\{|p_{\pi(i),n}^{k}-p_{i}|>c_{k}\delta\right\}\right)\leq\Pr\left(\bigcup_{i=1}^{k}\{H_{n}\left(B(\mathbf{a}_{i},\epsilon)\right)<p_{i}-\delta\}\right),$
@@
$q_{1}=q_{1}(i,\epsilon,n):=\Pr\left({\mathbf{X}_{1}/\|\mathbf{X}_{1}\|_{(s)}\in B(\mathbf{a}_{i},\epsilon),\ \|\mathbf{X}_{1}\|_{(r)}\geq\left(n/\ell_{n}\right)^{1/\alpha}}\right)/\Pr\left(\|\mathbf{X}_{1}\|_{(r)}\geq\left(n/\ell_{n}\right)^{1/\alpha}\right)\rightarrow p_{i}$ (36)
as $n\rightarrow\infty$, where the last convergence holds due to (8) and the fact that $B(\mathbf{a}_{i},\epsilon)$’s are disjoint under $\epsilon<r_{A}$, and
@@
$\Pr\left(\bigcup_{i=1}^{k}\{H_{n}\left(B(\mathbf{a}_{i},\epsilon)\right)<p_{i}-\delta\}\right)$ $\leq\sum_{i=1}^{k}\Pr\left(H_{n}(B(\mathbf{a}_{i},\epsilon))<p_{i}-\delta\right)$
@@
The large-deviation-type estimates in Proposition 4.3 say that the probability $\Pr\left(E(x,y)\right)$ decays exponentially in the expected extremal subsample size $c_{(r)}\ell_{n}$. It is worth observing that the expression of $\Delta(x,y)$ reflects the following: The difficulty of clustering-based estimation measured by the aforementioned large error probabilities depends negatively on $p_{min}$ and positively on $k$.
@@
Under the assumption and notation of Theorem 3.1, fix $t\in(0,t_{0})$. Then
$\limsup_{n}\frac{1}{c_{(r)}\ell_{n}}\ln\left\{\Pr\left(S_{t}(W_{n};A_{k,n},\mathfrak{C}_{k,n})\leq S_{t}(W_{n};A_{m,n},\mathfrak{C}_{m,n})\text{ for all }m\neq k\right)\right\}\leq\exp\left(-2\delta_{t}(k,p_{\min})^{2}\right)-1$
where $\delta_{t}(k,p_{\min})>0$ is the solution $\delta$ of the equation $[k(p_{\min}-\delta)r_{A}]^{t}-k\delta=\left(k^{2}\delta\right)^{t}\vee\left(\left(1-\left(p_{\min}-\delta\right)r_{A}\right)\mathbf{1}\left\{k\geq 2\right\}\right)$.
@@
and $\epsilon$ is sufficiently small (depending on $\delta$). Note that the inequality above holds when $\delta$ is sufficiently small due to $0<t<t_{0}=\ln\left(1-r_{A}p_{\min}\right)/\ln\left(r_{A}kp_{\min}\right)$, and its left-hand side is decreasing (to negative values) and its right-hand side is increasing with as $\delta$ increases to $p_{\min}$. Then for any $\delta\in(0,\delta_{t}(k,p_{\min}))$, we have in view of (37) and (4) that
@@
The proof is concluded by letting $\delta\uparrow\delta_{t}(k,p_{\min})$. ∎
@@ 5.1 The models
As observed by Einmahl et al. [5] and Janßen and Wan [20], one may relate a $k$-clustering algorithm to the estimation of certain factor-like models that are often considered in the analysis of multivariate extremes. Suppose $B=\left(b_{ij}\right)_{i=1,\ldots,d,j=1,\ldots,k}=\left(\mathbf{b}_{1},\ldots,\mathbf{b}_{k}\right)$, where $\mathbf{b}_{j}=(b_{1j},\ldots,b_{dj})^{\top}$, $j\in\{1,\ldots,k\}$, are $k$ distinct $d$-dimensional vectors, $b_{ij}\geq 0$, and that each column and row vector of $B$ is nonzero (otherwise, the dimension $d$ or the factor order $k$ can be reduced). Assume that $\mathbf{Z}=(Z_{1},\ldots,Z_{k})^{\top}$ is a vector of i.i.d.  positive continuous random variables satisfying $\Pr(Z_{1}>z)\sim z^{-\alpha}$ as $z\rightarrow\infty$, $\alpha\in(0,\infty)$. Then the sum-linear model is given as
@@
It is known that both models above satisfy MRV with (1), and have a discrete spectral measure as in (18) with
@@
This can be derived based on the well-known “single large jump” heuristic: when $\|\mathbf{X}\|_{(r)}$ is large, it is only due to a single large $Z_{j}$ with overwhelming probability. See, e.g., [22] and [5]; we mention that these works usually assume the same norm $\|\cdot\|_{(r)}=\|\cdot\|_{(s)}$ and $\alpha=1$, although an extension is straightforward. In addition, the marginal standardization condition (1) or equivalently (7) imposes the following restriction on $B$:
@@ 6.1 Simulation studies
In this section, we present some simulation studies to illustrate the performance of the penalized ASW method introduced in Section 3. We follow the setup in [20, Section 4] to simulate the max-linear factor model (40) with randomly generated coefficient matrix $B$. In particular, we let the factors $Z_{j}$’s each follow a standard Fréchet ($\alpha=1$) distribution. We consider 4 different combinations of dimensionality $d$ and true order $k$. Under each $(d,k)$ combination, we describe in the list below the way the coefficient vector $\mathbf{b}_{j}$’s are generated. Note that due to the standardization (42), only $\mathbf{b}_{1},\ldots,\mathbf{b}_{k-1}$ need to be specified. Let $U_{i}$’s stand for i.i.d. uniform random variables on $[0,1]$.
@@
For each of the 4 simulation setups described above, we randomly generate 100 models (i.e, 100 coefficient $B$ matrices). From each of these generated models, we simulate a dataset of size 1000, extract a subsample of size $100$ with the largest $2$-norms, and project the subsample on the $2$-norm sphere, namely, we work with $\|\cdot\|_{(r)}=\|\cdot\|_{(s)}=\|\cdot\|_{2}$. Subsequently, a spherical clustering algorithm (spherical $k$-means or $k$-pc) and the computation of the penalized ASW score is carried out on this projected subsample. Throughout the paper, for the spherical $k$-means algorithm, we use the implementation in the R package skmeans [16], and for the $k$-pc algorithm, we use the R implementation provided in the supplementary material of [7].
@@
In all these simulation setups, we can observe a tendency for the non-penalized ($t=0$) ASW to overestimate (sometimes greatly) the order. As the penalty parameter $t$ is tuned up from 00, we observe a significant bias correction effect, and the order identification success rate is noticeably improved over a range of $t>0$. Note that this success rate is calculated with respect to the same $t$ for different simulated data sets. We expect the success rate to improve if $t$ is adaptively tuned for each dataset following the visual method described in Section 3. It is also worth mentioning that the order identification based on $k$-pc tends to be more accurate than that based on $k$-means in most of these simulations.

### v2

@@ 3.2.1 Some deterministic estimates
Note that $\epsilon^{\prime}\to 0$ as $\epsilon,\delta\rightarrow 0$, we may choose them small enough such that (31) is satisfied. Define the event
$E_{n}(\epsilon,\delta)=\{|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|\geq|W_{n}|(p_{i}-\delta)\text{ for all }i\in\left\{1,\ldots,k\right\}\}.$ (34)
By Proposition 2.5 with $S=B_{D}(\mathbf{a}_{i},\epsilon)$ and the choice $\epsilon<r_{A}$, we have each $H_{n}(S)=|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|/|W_{n}|$ converges almost surely to $H(S)=p_{i}$, $i\in\{1,\ldots,k\}$. Hence, with probability $1$, the event $E_{n}(\epsilon,\delta)$ happens eventually as $n\rightarrow\infty$, namely, $\Pr\left(\liminf_{n}\mathbf{1}\left\{E_{n}(\epsilon,\delta)\right\}=1\right)=1$. Since $W_{n}$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$ on $E_{n}(\epsilon,\delta)$, by Lemmas 3.3 and 3.6, for almost every outcome $\omega$ in the sample space $\Omega$, when $n$ is sufficiently large, we have when $m<k$ that
@@
The desired results follow if one takes $\limsup_{n}$ and $\liminf_{n}$ respectively in the two inequalities above, and then lets $\delta,\epsilon\rightarrow 0$ (see also Remark 3.2). ∎
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Suppose $t>0$. If $m>k$, we have almost surely
@@
$\limsup_{n}P_{t}(W_{n};A_{k,n},\mathfrak{C}_{k,n})\leq 1-\left(r_{A}kp_{\min}\right)^{t},$
where $r_{A}$ is as in (20) and $p_{\min}$ is as in (24).
@@
Putting together Propositions 3.7 and 3.8, and using the facts that $\bar{S}\in[0,1]$ and $P_{t}\in[0,1]$, we have almost surely that
@@ 4 Large deviation analysis of clustering-based spectral estimation
First, we prepare a Chernoff-Hoeffding-type bound for the sum of a Binomial random number of Bernoulli random variables, which may be of some independent interest.
@@
We only prove the (35) and the proof of (36) is similar. It follows from a version of Hoeffding’s inequality for Binomial [18, Equation (2.1)] that for any $m\geq 0$,
@@
where in the last inequality we have used the inequality $x+1\leq\exp\left(x\right)$, $x\in\mathbb{R}$. To obtain the second inequality in (35), it suffices to note that in view of [18, Equation (2.3)] one has $\mathcal{D}(q_{1}-r|q_{1})\geq 2r^{2}$. ∎
@@
Let $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$ be as defined in (18). Let $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of the extremal subsample $W_{n}$ as in (14). By Corollary 2.6, there exists permutation $\pi$, such that $\mathbf{a}_{\pi_{n}(i),n}^{k}$ and $p_{\pi_{n}(i),n}^{k}$ in (19) are consistent estimators for $\mathbf{a}_{i}$ and $p_{i}$, respectively. Note that an accurate estimation can be interpreted as that for small $x,y>0$, $D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})<x$ and $|p_{\pi(i),n}^{k}-p_{i}|<y$ for all $i\in\{1,\ldots,k\}$. Now consider the complement “large deviation” event
$E(x,y)=\bigcap_{\pi}\bigcup_{i=1}^{k}\left\{|\mathbf{a}_{\pi(i),n}^{k}-\mathbf{a}_{i}|>x\right\}\cup\left\{|p_{\pi(i),n}^{k}-p_{i}|>y\right\}.$ (37)
where the intersection $\cap_{\pi}$ is over all permutations $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$. We have the following result.
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a spectral measure of the following form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form an $k$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Let $E(x,y)$ be the event defined in 37. Then for any $x,y>0$,
@@
If $H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\right)=|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|/|W_{n}|\geq p_{i}-\delta$ for all $i\in\{1,\ldots,k\}$, by Lemmas 3.4 and 3.6, as long as (31) holds, there exists a permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})<\epsilon^{\prime}$ and $|p_{\pi(i),n}^{k}-p_{i}|\leq c_{k}\delta$ for all $i\in\{1,\ldots,k\}$. Hence under (31), whenever $\epsilon^{\prime}\leq x$ or $c_{k}\delta\leq y$,
$\Pr\left(E(x,y)\right)\leq\Pr\left(\bigcap_{\pi}\bigcup_{i=1}^{k}\left\{D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})>\epsilon^{\prime}\right\}\cup\left\{|p_{\pi(i),n}^{k}-p_{i}|>c_{k}\delta\right\}\right)\leq\Pr\left(\bigcup_{i=1}^{k}\{H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\right)<p_{i}-\delta\}\right),$
@@
$q_{1}=q_{1}(i,\epsilon,n):=\Pr\left({\mathbf{X}_{1}/\|\mathbf{X}_{1}\|_{(s)}\in B_{D}(\mathbf{a}_{i},\epsilon),\ \|\mathbf{X}_{1}\|_{(r)}\geq\left(n/\ell_{n}\right)^{1/\alpha}}\right)/\Pr\left(\|\mathbf{X}_{1}\|_{(r)}\geq\left(n/\ell_{n}\right)^{1/\alpha}\right)\rightarrow p_{i}$ (38)
as $n\rightarrow\infty$, where the last convergence holds due to (8) and the fact that $B_{D}(\mathbf{a}_{i},\epsilon)$’s are disjoint under $\epsilon<r_{A}$, and
@@
$\Pr\left(\bigcup_{i=1}^{k}\{H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\right)<p_{i}-\delta\}\right)$ $\leq\sum_{i=1}^{k}\Pr\left(H_{n}(B_{D}(\mathbf{a}_{i},\epsilon))<p_{i}-\delta\right)$
@@
The large-deviation-type estimates in Proposition 4.3 say that the probability $\Pr\left(E(x,y)\right)$ decays exponentially in the expected extremal subsample size $c_{(r)}\ell_{n}$. It is worth observing that the expression of $\Delta(x,y)$ reflects the following: The difficulty of clustering-based estimation measured by the aforementioned large error probabilities depends negatively on $p_{min}$ and $r_{A}$ (note that $\epsilon_{0}$ is an increasing function of $r_{A}$) and positively on $k$. In other words, the estimation is more accurate when the true discrete spectral measure has fewer atoms, the dissimilarity between the atoms is large, or the probability mass concentrated on each atom is significant.
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Let $r_{A}$ be defined as in (20), $p_{\min}$ be defined as in (24) and $t_{0}$ be defined as in (25). Then fix $t\in(0,t_{0})$,
$\limsup_{n}\frac{1}{c_{(r)}\ell_{n}}\ln\left\{\Pr\left(S_{t}(W_{n};A_{k,n},\mathfrak{C}_{k,n})\leq S_{t}(W_{n};A_{m,n},\mathfrak{C}_{m,n})\text{ for all }m\neq k\right)\right\}\leq\exp\left(-2\delta_{t}(k,p_{\min},r_{A})^{2}\right)-1$
where $\delta_{t}(k,p_{\min},r_{A})>0$ is the solution $\delta$ of the equation $[k(p_{\min}-\delta)r_{A}]^{t}-k\delta=\left(k^{2}\delta\right)^{t}\vee\left(\left(1-\left(p_{\min}-\delta\right)r_{A}\right)\mathbf{1}\left\{k\geq 2\right\}\right)$.
@@
and $\epsilon$ is sufficiently small (depending on $\delta$). Note that the inequality above holds when $\delta$ is sufficiently small due to $0<t<t_{0}=\ln\left(1-r_{A}p_{\min}\right)/\ln\left(r_{A}kp_{\min}\right)$, and its left-hand side is decreasing (to negative values) and its right-hand side is increasing with as $\delta$ increases to $p_{\min}$. Then for any $\delta\in(0,\delta_{t}(k,p_{\min},r_{A}))$, we have in view of (39) and (4) that
@@
The proof is concluded by letting $\delta\uparrow\delta_{t}(k,p_{\min},r_{A})$. ∎
@@ 5.1 The models
As observed by Einmahl et al. [7] and Janßen and Wan [23], one may relate a $k$-clustering algorithm to the estimation of certain factor-like models that are often considered in the analysis of multivariate extremes. Suppose $B=\left(b_{ij}\right)_{i=1,\ldots,d,j=1,\ldots,k}=\left(\mathbf{b}_{1},\ldots,\mathbf{b}_{k}\right)$, where $\mathbf{b}_{j}=(b_{1j},\ldots,b_{dj})^{\top}$, $j\in\{1,\ldots,k\}$, are $k$ distinct $d$-dimensional vectors, $b_{ij}\geq 0$, and that each column and row vector of $B$ is nonzero (otherwise, the dimension $d$ or the factor order $k$ can be reduced). Assume that $\mathbf{Z}=(Z_{1},\ldots,Z_{k})^{\top}$ is a vector of i.i.d.  positive random variables satisfying $\Pr(Z_{1}>z)\sim z^{-\alpha}$ as $z\rightarrow\infty$, $\alpha\in(0,\infty)$. Then the sum-linear model is given as
@@
It is known that both models above satisfy MRV (2), and have a discrete spectral measure as in (18) with
@@
This can be derived based on the well-known “single large jump” heuristic: when $\|\mathbf{X}\|_{(r)}$ is large, it is only due to a single large $Z_{j}$ with overwhelming probability. See, e.g., [26] and [7]; we mention that these works usually assume the same norm $\|\cdot\|_{(r)}=\|\cdot\|_{(s)}$ and $\alpha=1$, although an extension is straightforward. In addition, the marginal standardization condition (1) or equivalently (7), imposes the following restriction on $B$:
@@ 6.1 Simulation studies
In this section, we present some simulation studies to illustrate the performance of the penalized ASW method introduced in Section 3. We follow the setup in [23, Section 4] to simulate the max-linear factor model (42) with randomly generated coefficient matrix $B$. In particular, we let the factors $Z_{j}$’s each follow a standard Fréchet ($\alpha=1$) distribution. We consider 4 different combinations of dimensionality $d$ and true order $k$. Under each $(d,k)$ combination, we describe in the list below the way the coefficient vector $\mathbf{b}_{j}$’s are generated. Note that due to the standardization (44), only $\mathbf{b}_{1},\ldots,\mathbf{b}_{k-1}$ need to be specified. Let $U_{i}$’s stand for i.i.d. uniform random variables on $[0,1]$.
@@
For each of the 4 simulation setups described above, we randomly generate 100 models (i.e, 100 coefficient $B$ matrices). From each of these generated models, we simulate a dataset of size $10000$, extract a subsample of size $1000$ with the largest $2$-norms, and project the subsample on the $2$-norm sphere, namely, we work with $\|\cdot\|_{(r)}=\|\cdot\|_{(s)}=\|\cdot\|_{2}$. Subsequently, a spherical clustering algorithm (spherical $k$-means or $k$-pc) and the computation of the penalized ASW score is carried out on this projected subsample. Throughout the paper, for the spherical $k$-means algorithm, we use the implementation in the R package skmeans [19], and for the $k$-pc algorithm, we use the R implementation provided in the supplementary material of [10].
@@
In all these simulation setups, we can observe a tendency for the non-penalized ($t=0$) ASW to overestimate (sometimes greatly) the order. As the penalty parameter $t$ is tuned up from 00, we observe a significant bias correction effect, and the order identification success rate is noticeably improved over a range of $t>0$. Note that this success rate is calculated with respect to the same $t$ for different simulated data sets. We expect the success rate to improve if $t$ is adaptively tuned for each dataset following the visual method described in Section 3. It is also worth mentioning that the order identification based on $k$-pc tends to be more accurate than that based on $k$-means in most of these simulations. Moreover, the lower success rate for $d=4,k=6$ seems to arise because this configuration tends to produce centers that are in proximity to one another.
In the supplementary material [5], we include more simulation studies as above with different choices of extremal subsample sizes. The results are similar in terms of the bias correction effect of the penalty. We also observe that when the extremal subsample size is chosen relatively large (e.g. $1500$), the accuracy of order selection deteriorates, which is expected since the effect of the convergence of Corollary 2.6 is not yet strong with a low threshold.